OpenStudy (anonymous):

5. What is the approximate radius of a sphere whose volume is 1,349 cm³?

OpenStudy (anonymous):

The volume of a sphere is \[(4/3)\pi r ^{3}\] Plugging in we get: \[1349=(4/3)\pi r ^{3}\] Now you just have to solve.

OpenStudy (anonymous):

so i just solve that? or do i need to get the radius first?

OpenStudy (anonymous):

Yes, just solve. You are solving for the radius aren't you? The radius will be your only unknown, so you don't need it.

OpenStudy (anonymous):

oh allright...thanks so much..so when i get to 322.1 = r^3 i just find the cube of 322.1?

OpenStudy (anonymous):

Cube root

OpenStudy (anonymous):

how do you do a cube root?

OpenStudy (anonymous):

on a calculator

OpenStudy (anonymous):

\[\sqrt[3]{322.1}\] What type of calculator are you using?

OpenStudy (anonymous):

just a downloaded one...ill try tht

OpenStudy (anonymous):

i cant just do the square root twice?

OpenStudy (anonymous):

No, that won't work. Hold on I'll get it.

OpenStudy (anonymous):

Here you go... http://www.wolframalpha.com/input/?i=cube%20root%20of%20322.1&t=crmtb01

OpenStudy (anonymous):

i figured it out...thanks so much